OpenStudy (anonymous):

Will give medal. In five years, a garden festival plans on using the park where Iris has been studying the invasive snails. Explain to the garden festival committee how to find the average rate of change for the snail population between years 2 and 5. Describe what this average rate of change represents. For year 2 there where 40 snails Year 5 there where 360 snails. I got 93.3333333 I don't know what this means.

OpenStudy (anonymous):

you basically finding the slope. do you remember the slope formula. you just do y2-y2/x2-x1. Also Time is always the x axis. so snail pop will be the y-axis.

OpenStudy (anonymous):

ok. That's how I ended up with 93.333\[\frac{ 360-40 }{ 5-2 } = \frac{ 280 }{ 3 } = 93.33333\]

OpenStudy (anonymous):

360-40 is not 280

OpenStudy (anonymous):

sorry the 360 was supposed to be a 320. My bad

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

The 93.3 means that the on average there are 93 snails/year.

OpenStudy (anonymous):

do you understand?

OpenStudy (anonymous):

I understand it just doesn't make sense with the function or table. The function is \[f(x)=10(2)^{x}\]

OpenStudy (anonymous):